# ICYM2I: The illusion of multimodal informativeness under missingness

### v2

@@ Abstract
Multimodal learning is of continued interest in artificial intelligence-based applications, motivated by the potential information gain from combining different types of data. However, modalities observed in the source environment may differ from the modalities observed in the target environment due to multiple factors, including cost, hardware failure, or the perceived informativeness of a given modality. This shift in missingness between the source and target environment has not been carefully studied. Naïve estimation of the information gain associated with including an additional modality without accounting for missingness may result in improper estimates of that modality’s value in the target environment. We formalize the problem of missingness, demonstrate its ubiquity, and show that the subsequent distribution shift results in bias when the missingness process is not explicitly accounted for. To address this issue, we introduce ICYM2I (In Case You Multimodal Missed It), a framework [^1] Available on Github https://github.com/reAIM-Lab/ICYM2I/ for the evaluation of predictive performance and information gain under missingness through inverse probability weighting-based correction. We demonstrate the importance of the proposed adjustment to estimate information gain under missingness on synthetic, semi-synthetic, and real-world datasets.
@@ 1 Introduction
Multimodal learning is ubiquitous in machine learning as practitioners combine multiple data types to improve predictive performance in applications to healthcare [1, 2], robotics [3, 4], and recommender systems [5]. However, factors such as privacy concerns [6, 7], cost-benefit tradeoffs of data-acquisition [8], and user preferences [9] imprint multimodal data with missingness. Additionally, even if modality complete data is available or curated at training, data noise [10, 11] and sensor failures [12, 13] may result in missing modalities in the target environment.
Despite the existence of missingness in real-world settings, current multimodal machine learning methods often assume modalities are fully observed, both in source and target environments. When missingness is considered, the literature has focused on engineering efforts [14, 15] such as data selection [16], imputation [17, 18, 19], and architecture design [20, 21], which implicitly assume a stable missingness process between source and target environments. When this assumption is violated, the missingness mechanism induces a distribution shift [22, 23] that biases the estimated informativeness of a given modality. Missingness is pervasive and impacts a broad range of application domains encountered in the multimodal literature: in breast cancer screening, biopsies are only performed if there are abnormal findings in a mammogram; in autonomous vehicles, LiDAR sensor dropout can occur due to weather and lighting conditions; and in online recommender systems, reviews are only collected after certain consumer behaviors. Across these settings, ignoring the distribution shift between source and target due to missingness when quantifying modality informativeness may conflate missingness with signal, leading to flawed data collection and modeling decisions.
@@
  • ICYM2I. Under the missingness-at-random (MAR) assumption, a much more realistic assumption than the common and often implicit assumption of missingness-completely-at-random (MCAR) made by state-of-the-art multimodal strategies, we propose ICYM2I  (In Case You Multimodal Missed It), a double inverse-propensity weighting correction to overcome missingness-induced distribution shifts. Specifically, we demonstrate that ICYM2I improves correlation in predictive and information-theoretic utility of modalities.
@@ 2 Related Work
Multimodal benchmarks suppress missingness encountered in real-world environments. Prior work on multimodal models often assumes fully observed modalities [24, 25, 26]. Missingness has largely been an overlooked problem [14, 15], to the extent that current benchmarks rarely contain samples with missing modalities. Curation often involves dropping incomplete or filtering samples based on data quality criteria, such as text length or file size [27, 28] or imputing with automatic tools [29]. This curation implicitly assumes that rejected samples follow the same distribution as the observed ones. This assumption may not hold. For example, in autonomous driving data, samples with sensor failure – often resulting from extreme weather or lighting conditions – may be filtered out. Models trained on complete data may, consequently, not generalize to these settings, creating real-world risk at deployment. When missingness is considered, previous works focused on robustness through imputation [17, 18, 19], representation learning [15, 30], knowledge distillation [31, 32], and model ensembling [20, 21] – all ignoring the potential shift resulting from the missingness process.
@@
Multimodal missingness in the target environment. Prior work has explored missingness in the target environment [33, 21], e.g., when a captor fails at deployment [34]. Broadly, two strategies have been proposed [15]: (i) data preprocessing through cross-modal imputation [18, 19, 17], where one replaces the missing modality [35, 36], as well as (ii) model training strategies such as architecture design [37, 38], distillation-based methods [31, 32], and ensembling [20, 21]. Through the proposed formalization, our work distinguishes between different missingness assumptions, demonstrating that the previously studied framework is only one among various plausible mechanisms for which current strategies are not well-designed.
@@
Distribution shifts in multimodal learning. Addressing multimodal shifts has been studied in vision-language models [39, 40] or using information-theoretic notions to understand multimodal behavior under distribution shifts [41]. Augmentation and regularization strategies have been leveraged to address temporal shifts for conversation understanding [42, 43]. Advances in learning, such as in-context learning, have been studied to characterize adaptation to multimodal distribution shifts [39, 44]. However, existing strategies aim to improve robustness under domain distribution shifts, while ignoring the potential shift in missingness between source and target environments.
@@
Quantifying information-theoretic value of a modality. Existing works often implicitly assume that additional modalities improve performance, ignoring the prohibitive cost, complexity, and potential noise added by these additional dimensions. When limited resources or constraints limit availability in the target environment, a central challenge is to quantify the information-theoretic value of a modality [45]. Liang et al. [46] proposed a method for recovering partial information decomposition measures of the redundancy, uniqueness, and synergy of the information provided by the different modalities [47, 48]. However, these decompositions fully ignore the impact of missingness.
@@
Correcting for missingness bias. The lack of formalization of missingness in the multimodal literature has led to neglecting its potential impact. Ignoring this process risks biasing estimates of interest [49] as the observed distribution differs from the underlying one practitioners aim to model. The statistical literature has introduced strategies such as matching [50] and reweighting [51] to correct for the missingness process. However, these works have overlooked the multimodal setting and the systematic shifts that may occur in this setting.
@@ 3 Multimodality and missingness
Consider two modalities, $X_{1}\in\mathcal{X}_{1}$ and $X_{2}\in\mathcal{X}_{2}$ and the state of interest $Y\in\mathcal{Y}$. We denote the joint underlying distribution $\Omega=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{Y}$. Without loss of generality, we assume an anti-causal setting for the data-generating process, in which the modalities are dependent on the states $Y$, as shown in Figure 2 (left). We use the binary indicators of missingness $M_{1}$, $M_{2}$, and $M_{Y}$, which are equal to $1$ if the associated variable is missing, 0 if observed, following the convention of Mohan and Pearl [52]. Observed variables are subscripted by ‘obs’, which corresponds to the underlying modality if observed, and unobserved otherwise (denoted by $\varnothing$). Formally, the observed variable $Y_{obs}$ and observed modalities $X_{1,obs}$ and $X_{2,obs}$ can be defined as follows:
@@
In this setting, we denote the observed joint distribution $\Omega_{\text{obs}}^{\text{}}:=(M_{1}\cdot\mathcal{X}_{1})\times(M_{2}\cdot\mathcal{X}_{2})\times(M_{Y}\cdot\mathcal{Y})$, which has been the focus of multimodal learning. Note that the underlying relation between the covariates of interest $X_{1}$, $X_{2}$, and $Y$ remains unchanged under the true data-generating mechanism; only the realizations of the observed variables differ.
Missingness in multimodal learning. We distinguish three mechanisms that cover potential missingness in multimodal data [53]:
@@
Missingness-induced distribution shifts. Missingness in a modality $X_{i}$ and/or the label $Y$ can induce distribution shifts between the source and the target distributions. For example, if a modality is observed in the target environment only if another one meets some criterion, then this distribution may not match the source distribution. Theoretically, we know that a non-MCAR missingness mechanism induces distribution shifts [23, 22], i.e., the observed distribution differs from the underlying distribution. Critically, models trained and evaluated on the observed distribution are statistically biased estimates under any other missingness process.
For instance, consider an autonomous vehicle setting where video and LiDAR represent two modalities of interest. If the LiDAR randomly dysfunctions, the missingness patterns are MCAR. However, as previously mentioned, LiDAR may malfunction under extreme weather conditions. If these conditions can be extracted from the video modality, one may assume MAR patterns. However, if the video cannot capture the variable explaining the LiDAR dysfunction – e.g., temperature – LiDAR would be MNAR, as dependent upon an unobserved variable. Under the last two scenarios, focusing solely on samples with LiDAR excludes all extreme condition settings.
A common and often implicit assumption in the multimodal literature is the absence of missingness, which is analogous to either a MCAR mechanism or a stable missingness process as described in Assumption A. In other words, not adjusting for missingness assumes that the missingness process is uninformative or will remain the same in the target environment.
Assumption A (Stable missingness process).
The missingness process is stable between the source and target environments, i.e., the observed distributions are the same $\Omega_{\text{obs}}^{\text{source}}=\Omega_{\text{obs}}^{\text{target}}$.
A few prior works focus on improving the robustness of multimodal models when performance may degrade due to a modality missing in the target distribution, as formalized in Assumption B.
Assumption B (Collection degradation).
Collection in the target environment may deteriorate, resulting in a distribution shift $\Omega_{\text{obs}}^{\text{source}}\sim\Omega^{\text{}}\neq\Omega_{\text{obs}}^{\text{target}}$.
Our work questions the applicability of these assumptions where modality collection is costly. While missingness may naturally present a distributional shift [54], we emphasize that demonstrating the value of a modality in the source environment may lead to increased collection of this modality in the target environment, inducing a distribution shift akin to Assumption C.
Assumption C (Multimodal analysis informs data collection).
@@
We focus on settings where historical data on which one can train a model is marked by missingness. Under such settings, we aim to do the following: (i) identify which modalities are informative and may consequently be fully observed in the target environment, and (ii) train models that generalize to the full observation of these modalities.
@@ 4 Is this modality informative?
We aim to assess whether a partially missing modality would be informative if fully observed. To this end, we introduce ICYM2I (In Case You Multimodal Missed It), a framework for correcting model performance trained on modality complete samples where all modalities and labels are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$ to estimate the predictive utility of the partially missing modality if it were observed for the whole population ($\Omega^{\text{}}$). Additionally, we propose a correction to derive unbiased estimates of the information-theoretic utility of a modality, using $\Omega_{\text{obs}}^{\text{}}$. We rely on Partial Information Decomposition (PID) [48] bounds introduced by Bertschinger et al. [47] for this task, which quantifies the information value of a target of interest captured by two input variables.
Correction. We propose an Inverse Probability Weighting (IPW) variant [55], which reweights samples based on their probability of being observed. Under Assumption D, IPW recovers unbiased estimates of the true distribution, enabling learning and evaluation on the true distribution from observed samples. IPW-adjustment is critical for both training and evaluation of multimodal models under missingness. IPW-adjusted training results in a model trained to infer on the underlying distribution $\Omega^{\text{}}$, while correction of the evaluation allows for measuring performance on $\Omega^{\text{}}$, despite evaluating the model only on samples from the observed distribution $\Omega_{\text{obs}}^{\text{}}$.
Assumption D (MAR and Positivity).
@@ 4.1 A motivating example
We consider the common multimodal example of learning bit-wise logic operators [47, 56, 46]. We generate $10,000$ points with two modalities drawn from Bernoulli distributions ($p=0.5$). The output state $Y$ is defined using the binary operators AND, OR, and XOR of input bits $X_{1}$ and $X_{2}$. In this setting, we induce missingness $M_{2}$ in $X_{2}$ and $Y$ as a function of $X_{1}$ (MAR): $M_{2}\sim Bern(0.6X_{1}+0.2)$, resulting in 50% missingness in $X_{2}$. We investigate the impact of missingness on current strategies for evaluating the predictive and information-theoretic utility of a given modality.
Estimating performance for informativeness. A common practice to measure the predictive value of adding a modality is through modality ablation studies where practitioners train models on the subset of observed samples where all modalities are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$. First, unimodal models $f_{\phi}(x_{i})$ are trained to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$ and a multimodal model $f_{\phi}(x_{1},x_{2})$ to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{1},x_{2})$ on the same observed dataset. The performance is then compared in a hold-out set, which is also sampled from $\Omega_{\text{obs}}^{\text{}}$. The informativeness of a modality is heuristically attributed to the relative performance gain of the multimodal model compared to the unimodal model. However, multimodal models can perform worse than their unimodal counterparts due to data characteristics [57] and learning dynamics [58, 59]. Thus, relying solely on performance as a proxy for informativeness, particularly under distribution shifts, can be misleading.
Partial Information Decomposition [47]. As an alternative to estimating performance, existing works have decomposed the informativeness associated with each modality [46]. Partial information decomposition (PID) [47, 48] decomposes the total mutual information $I(Y:(X_{1},X_{2}))$ [60, 61] between a random variable $Y$ and $(X_{1}$, $X_{2})$ jointly into shared information (information both $X_{1}$, $X_{2}$ share about $Y$), unique information 1 (information only $X_{1}$ has about $Y$), unique information 2 (information only $X_{2}$ has about $Y$), and complementary information (information about $Y$ that requires both $X_{1}$ and $X_{2}$):
@@
Let $\Delta$ be the space of all distributions over $(X_{1},X_{2},Y)$ and let $\Omega^{\text{}}$ denote the true data distribution (without missingness) and define $\Delta_{\Omega^{\text{}}}:=\big\{q\in{\Delta}:q(X_{i}=x_{i},Y=y)=p_{\Omega}^{\text{}}(X_{i}=x_{i},Y=y)\text{ }\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\big\}$. That is, $\Delta_{\Omega^{\text{}}}$ is the set of all distributions over $(X_{1},X_{2},Y)$ such that the two-way joints between $X_{i}$ and $Y$ match the true data-generating distribution. Equipped with this set, Bertschinger et al. [47] provides the following bounds $\widetilde{SI}$, $\widetilde{UI}$, and $\widetilde{CI}$ on the analogous quantities:
$\widetilde{SI}(Y:X_{1};X_{2})$ $:=\max_{Q\in{\Delta_{\Omega^{\text{}}}}}[I_{q}(Y:X_{1})-I_{q}(Y:X_{1}|X_{2})]$ $\leq SI(Y:X_{1};X_{2})$ (1)
@@
[…]
$\widetilde{CI}(Y:X_{1};X_{2})$ $:=I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))-\min_{Q\in{\Delta_{\Omega^{\text{}}}}}I_{q}(Y:(X_{1},X_{2}))$ $\leq CI(Y:X_{1};X_{2})$ (4)
These bounds are tight if there exists a $q_{0}\in\Delta_{\Omega^{\text{}}}$ such that $\widetilde{CI}_{q_{0}}(Y:X_{1};X_{2})=0$. Bertschinger et al. [47] further shows that under common (but unverifiable) assumptions on the data-generating process, the inequalities are tight for all $q\in\Delta_{\Omega^{\text{}}}$. This results in a compelling argument to rely on these entities, as it suggests that it is not possible to decide whether or not complementary information exists when only marginals $(Y,X_{1})$ and $(Y,X_{2})$ are known. Hence, we choose this measure of PID. Prior works relying on this PID to attribute information-theoretic value implicitly assume that $\Omega_{\text{obs}}^{\text{source}}=\Omega^{\text{source}}=\Omega^{\text{target}}=\Omega$. Instead, we evidence the limitations of these strategies performed on $\Omega_{\text{obs}}^{\text{source}}$ when the target decomposition is $\Omega^{\text{target}}=\Omega$, i.e., the true data-generating mechanism.
@@
As a motivating example, we analyze the impact of missingness on estimating PID in the case of unidimensional modalities with a bitwise logic outcome (AND, OR, and XOR). Table 1 (left) presents the discriminative performance associated with neural networks trained on each modality and their combination under three scenarios: (i) access to all data (Oracle), (ii) focusing only on datapoints with all covariates observed (Observed), and (iii) adequately accounting for missingness (ICYM2I  using IPW to adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$, by modeling the missingness mechanism), as proposed in Section 4.2. Table 1 (right) presents PID, discussed in Section 4.3 under the same scenarios, demonstrating how information decomposition is also biased due to missingness.
Specifically, relying on $\Omega_{\text{obs}}^{\text{}}$ overestimates the performance of $X_{1}$ for OR but underestimates it for AND. Similarly, biased decomposition results in overestimating the informativeness of $X_{1}$ (“Unique 1” compared to “Unique 2”) for OR. As $X_{1}$ informs the missingness process, it indirectly informs the outcome of interest, despite the true underlying generative process being dependent on both. The use of IPW can correct for such bias under positivity as long as the propensities for IPW can be estimated (i.e., the MAR assumption). We study sensitivity to this assumption in Appendix D, where we further evaluate the robustness of our method under MCAR and MNAR, demonstrating robustness under MCAR.

### v3

@@ Abstract
Multimodal learning is of continued interest in artificial intelligence-based applications, motivated by the potential information gain from combining different data modalities. However, modalities observed in the source environment may differ from the modalities observed in the target environment due to multiple factors, including cost, hardware failure, or the perceived informativeness of a given modality. This change in missingness patterns between the source and target environment has not been carefully studied. Naïve estimation of the information gain associated with including an additional modality without accounting for missingness may result in improper estimates of that modality’s value in the target environment. We formalize the problem of missingness, demonstrate its ubiquity, and show that the subsequent distribution shift induces bias when the missingness process is not explicitly accounted for. To address this issue, we introduce ICYM2I (In Case You Multimodal Missed It), a framework [^1] Code available on Github: https://github.com/reAIM-Lab/ICYM2I for the evaluation of predictive performance and information gain under missingness through inverse probability weighting-based correction. We demonstrate the importance of the proposed adjustment to estimate information gain under missingness on synthetic, semi-synthetic, and real-world datasets.
@@ 1 Introduction
Multimodal learning is ubiquitous in machine learning as practitioners combine multiple data types to improve predictive performance in applications to healthcare (Perochon et al., 2023; Tu et al., 2024), robotics (Gao et al., 2024; Shah et al., 2023), and recommender systems (Chen et al., 2019). However, factors such as privacy concerns (Jaiswal and Provost, 2020; Zhang et al., 2021), cost-benefit tradeoffs of data-acquisition (Buck et al., 2010), and user preferences (Kossinets, 2006) imprint multimodal data with missingness. Additionally, even if modality complete data is available or curated at training, data noise (Cohen et al., 2004; Ma et al., 2023) and sensor failures (Inceoglu et al., 2021; 2023) may result in missing modalities in the target environment.
Although the missingness of modalities is a recurring challenge in real-world settings, current multimodal machine learning methods often assume that modalities are fully observed, both in source and target environments. When missingness is considered, the literature has focused on engineering efforts (Le et al., 2025; Wu et al., 2024) such as data selection (Hosseini et al., 2022), imputation (Tran et al., 2017; Cohen Kalafut et al., 2023; Malitesta et al., 2024), and architecture design (Chen et al., 2022; Zeng et al., 2022), which implicitly assume a stable missingness process between source and target environments. When this assumption is violated, the missingness mechanism induces a distribution shift (Zhang et al., 2023; Liu et al., 2023b) that biases the estimated informativeness of a given modality. Missingness is pervasive and impacts a broad range of application domains encountered in the multimodal literature: in breast cancer screening, biopsies are only performed if there are abnormal findings in a mammogram; in autonomous vehicles, LiDAR sensor dropout can occur due to weather and lighting conditions; and in online recommender systems, reviews are only collected after certain consumer behaviors. Across these settings, ignoring the distribution shift between source and target due to missingness when quantifying modality informativeness may conflate missingness with signal, leading to flawed data collection and modeling decisions.
@@
  • ICYM2I. Under the missingness-at-random (MAR) assumption, a much more realistic assumption than the common and often implicit assumption of missingness-completely-at-random (MCAR) made by state-of-the-art multimodal strategies, we propose ICYM2I  (In Case You Multimodal Missed It), a double inverse-propensity weighting correction to overcome missingness-induced distribution shifts. The proposed method does not aim to improve predictive performance; rather, it enables unbiased estimation of model performance and information gain across modalities, even under missingness. Specifically, we demonstrate that ICYM2I improves correlation in predictive and information-theoretic utility of modalities.
@@ 2 Related Work
Multimodal benchmarks suppress missingness encountered in real-world environments. Prior work on multimodal models often assumes fully observed modalities (Ngiam et al., 2011; Zadeh et al., 2017; Hou et al., 2019). Missingness has largely been an overlooked problem (Le et al., 2025; Wu et al., 2024), to the extent that current benchmarks rarely contain samples with missing modalities. Curation often involves discarding incomplete samples, filtering them based on data-quality criteria, such as text length or file size (Sharma et al., 2018; Schuhmann et al., 2022), or imputing them using automatic tools (Miech et al., 2019). This curation implicitly assumes that rejected samples follow the same distribution as the observed ones. This assumption may not hold. For example, in autonomous driving datasets, samples with sensor failures – often caused by extreme weather or lighting conditions – may be filtered out. Models trained on complete data may consequently not generalize to these settings, posing real-world risk at deployment.
When missingness is considered, previous works focused on robustness through imputation (Tran et al., 2017; Cohen Kalafut et al., 2023; Malitesta et al., 2024), representation learning (Wu et al., 2024; Liu et al., 2023a), knowledge distillation (Li et al., 2024; Wang et al., 2020a), and model ensembling (Chen et al., 2022; Zeng et al., 2022) – all ignoring the potential shift resulting from the missingness process.
@@
Multimodal missingness in the target environment. Prior work has explored missingness in the target environment (Lin and Hu, 2023; Zeng et al., 2022), e.g., when a captor fails at deployment (Ma et al., 2022). Broadly, two strategies have been proposed (Wu et al., 2024): (i) data preprocessing through cross-modal imputation (Cohen Kalafut et al., 2023; Malitesta et al., 2024; Tran et al., 2017), where one replaces the missing modality (Ma et al., 2021; Zhou et al., 2022), as well as (ii) model training strategies such as architecture design (Lee et al., 2023; Ge et al., 2023), distillation-based methods (Li et al., 2024; Wang et al., 2020a), and ensembling (Chen et al., 2022; Zeng et al., 2022). Through the proposed formalization, our work distinguishes between different missingness assumptions, demonstrating that the previously studied framework is only one among various plausible mechanisms for which current strategies are not well-designed.
@@
Distribution shifts in multimodal learning. Addressing multimodal shifts has been studied in vision-language models (Zhou et al., 2024; Verma et al., 2024) or using information-theoretic notions to understand multimodal behavior under distribution shifts (Oh et al., 2025). Augmentation and regularization strategies have been leveraged to address temporal shifts for conversation understanding (Woo et al., 2023; Lian et al., 2023). Advances in learning, such as in-context learning, have been studied to characterize adaptation to multimodal distribution shifts (Zhou et al., 2024; Xue et al., 2024). However, existing strategies aim to improve robustness under domain shifts only, while ignoring the potential shift in missingness between source and target environments. Instead, our work aims to correct estimates of performance and modality informativeness under missingness to inform modality collection at deployment.
@@
Quantifying information-theoretic value of a modality. Existing works often implicitly assume that additional modalities improve performance, ignoring the prohibitive cost, complexity, and potential noise added by these additional dimensions. When limited resources or constraints limit availability in the target environment, a central challenge is to quantify the information-theoretic value of a modality (Liang et al., 2024c). Liang et al. (2024a) proposed a method for recovering partial information decomposition measures of the redundancy, uniqueness, and synergy of the information provided by the different modalities (Bertschinger et al., 2014; Williams and Beer, 2010). However, these decompositions ignore the impact of missingness, a gap that our work aims to fill.
@@
Correcting for missingness bias. The impact of missingness in multimodal learning is understudied, in part due to the lack of formalization of missingness in this area. Ignoring this process can bias estimates of interest (Phelan et al., 2017), because the observed distribution differs from the underlying distribution that practitioners aim to model. The statistical literature has introduced strategies such as matching (Stuart, 2010) and reweighting (Jethani et al., 2022) to address missingness. We leverage these statistical tools to address the overlooked missingness in multimodal settings.
@@ 3 Multimodality and missingness
Consider two modalities, $X_{1}\in\mathcal{X}_{1}$ and $X_{2}\in\mathcal{X}_{2}$ and the state of interest $Y\in\mathcal{Y}$. We denote the joint underlying distribution $\Omega=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{Y}$. Without loss of generality, we assume an anti-causal setting for the data-generating process, in which the modalities are dependent on the states $Y$, as shown in Figure 2 (left). We use the binary indicators of missingness $M_{1}$, $M_{2}$, and $M_{Y}$, which are equal to $1$ if the associated variable is missing, 0 if observed, following the convention of Mohan and Pearl (2021). Observed variables are subscripted by ‘obs’, which corresponds to the underlying modality if observed, and unobserved otherwise (denoted by $\varnothing$). Formally, the observed variable $Y_{obs}$ and observed modalities $X_{1,obs}$ and $X_{2,obs}$ can be defined as follows:
@@
In this setting, we denote the observed joint distribution $\Omega_{\text{obs}}^{\text{}}:=(M\cdot\mathcal{X}_{1})\times(M\cdot\mathcal{X}_{2})\times(M\cdot\mathcal{Y})$ where $M=M_{1}\cdot M_{2}\cdot M_{Y}$. This complete modalities analysis has been the focus of multimodal learning.
Missingness in multimodal learning. We distinguish three mechanisms that cover potential missing modality in multimodal settings (Rubin, 1976):
@@
Missingness-induced distribution shifts. Missing modality $X_{i}$ and/or missing label $Y$ can induce distribution shifts between the source and the target distributions. For example, if a modality is observed in the target environment only if another one meets some criterion, then this distribution may not match the source distribution. Theoretically, we know that a non-MCAR missingness mechanism induces distribution shifts (Liu et al., 2023b; Zhang et al., 2023), i.e., the observed distribution differs from the underlying distribution. Critically, models trained and evaluated on the observed distribution are statistically biased estimates under any other missingness process.
For instance, consider an autonomous vehicle setting in which video and LiDAR are two modalities of interest. If the LiDAR sensor randomly fails, the missingness is MCAR. However, as previously mentioned, LiDAR may be more likely to malfunction under extreme weather conditions. If these conditions can be extracted from the video modality, one may assume MAR patterns. However, if the video cannot capture the variable that explains the LiDAR dysfunction (e.g., temperature), then LiDAR would be MNAR, as it is dependent on an unobserved variable. In the last two scenarios, focusing solely on complete samples would exclude all extreme-condition settings with critical real-world repercussions.
A common and often implicit assumption in the multimodal literature is the absence of missingness, which corresponds to either a MCAR mechanism ($\Omega_{\text{obs}}^{\text{}}=\Omega^{\text{}}$) or a stable missingness process between the source and target environments, i.e., the source and target distributions are the same ($\Omega_{\text{obs}}^{\text{source}}=\Omega_{\text{obs}}^{\text{target}}$). In other words, not adjusting for missingness assumes that the missingness process is uninformative or will remain the same in the target environment.
Prior works focus on improving the robustness of multimodal models when performance may degrade due to a modality missing in the target distribution, i.e, the source distribution reflects the true distribution while the target may present missingness $\Omega_{\text{obs}}^{\text{source}}\sim\Omega^{\text{}}\neq\Omega_{\text{obs}}^{\text{target}}$.
Our work questions the applicability of these assumptions where modality collection is costly. While missingness may result in a distributional shift (Zhou et al., 2023), we emphasize that demonstrating the value of a modality in the source environment may lead to increased collection of this modality in the target environment, inducing a distribution shift akin to Assumption A.
Assumption A (Multimodal analysis informs data collection).
@@
We focus on settings in which historical data used to train a model are marked by missingness. Under such settings, we aim to: (i) identify which modalities are informative and may consequently be collected in the target environment, and (ii) train models using the observed source data that generalize to the target environment where the modalities are fully observed.
@@ 4 Is this modality informative?
We aim to assess whether a partially missing modality would be informative if fully observed. To this end, we introduce ICYM2I (In Case You Multimodal Missed It), a framework for correcting model performance trained on modality complete samples where all modalities and labels are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$ to estimate the predictive utility of the partially missing modality if it were observed for the whole population ($\Omega^{\text{}}$). Additionally, we propose a correction to derive unbiased estimates of the information-theoretic utility of a modality, using $\Omega_{\text{obs}}^{\text{}}$. We rely on Partial Information Decomposition (PID) (Williams and Beer, 2010) bounds introduced by Bertschinger et al. (2014) for this task, which quantifies the information value of a target of interest captured by two input variables.
Correction. We propose an Inverse Probability Weighting (IPW) approach (Robins et al., 1994), which reweights samples according to their probability of being observed. Under Assumption B that relaxes the common MCAR assumption made in the multimodal literature, IPW recovers unbiased estimates of the true distribution, enabling learning and evaluation on the true distribution from observed samples. IPW-adjustment is critical for both training and evaluation of multimodal models under missingness. IPW-adjusted training results in a model trained to infer on the underlying distribution $\Omega^{\text{}}$, while correction of the evaluation allows for measuring performance on $\Omega^{\text{}}$, despite evaluating the model only on samples from the observed distribution $\Omega_{\text{obs}}^{\text{}}$.
Assumption B (MAR and Positivity).
@@ 4.1 A motivating example
We consider the common multimodal example of learning bit-wise logic operators (Bertschinger et al., 2014; Harder et al., 2013; Liang et al., 2024a). We generate $10,000$ points with two modalities drawn from Bernoulli distributions ($p=0.5$). The output state $Y$ is defined using the binary operators AND, OR, and XOR of input bits $X_{1}$ and $X_{2}$. In this setting, we induce missingness $M_{2}$ in $X_{2}$ and $Y$ as a function of $X_{1}$ (MAR): $M_{2}\sim Bern(0.6X_{1}+0.2)$, resulting in 50% missingness in $X_{2}$. We investigate the impact of missingness on current strategies for evaluating the predictive and information-theoretic utility of a given modality.
Estimating performance for informativeness. A common practice to measure the predictive value of adding a modality is through ablation studies where practitioners train models using different modalities on the subset of samples where all modalities are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$. First, unimodal models $f(x_{i})$ are trained to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$ and a multimodal model $f(x_{1},x_{2})$ to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{1},x_{2})$ on the same observed dataset. The performance is then compared in a hold-out set, sampled from $\Omega_{\text{obs}}^{\text{}}$. The informativeness of a modality is measured by the relative performance gain of the multimodal model over the unimodal model. However, multimodal models can underperform their unimodal counterparts due to data characteristics (Zhang et al., 2024) and learning dynamics (Wang et al., 2020b; Zhai et al., 2024). Thus, relying solely on performance as a proxy for informativeness, particularly under distribution shifts, can be misleading.
Partial Information Decomposition (Bertschinger et al., 2014). As an alternative to estimating performance, existing works have decomposed the informativeness associated with each modality (Liang et al., 2024a). Bertschinger et al. (2014) formalized this decomposition by analyzing the total (three-way) mutual information $I(Y:(X_{1},X_{2}))$ (McGill, 1954; Te Sun, 1980), a measure of dependency between the target variable $Y$ and the modalities $(X_{1}$, $X_{2})$. This quantity can be decomposes into four components: shared information (information both $X_{1}$, $X_{2}$ share about $Y$), unique information 1 (information only $X_{1}$ has about $Y$), unique information 2 (information only $X_{2}$ has about $Y$), and complementary information (information about $Y$ that requires both $X_{1}$ and $X_{2}$) as follows:
@@
Bertschinger et al. (2014) specifies how to estimate these quantities (see Appendix B); for instance, the unique information between $Y$ and $X_{1}$ can be estimated using the following:
@@
where $\Delta_{\Omega^{\text{}}}$ is the set $q(X_{i}=x_{i},Y=y)=p_{\Omega}^{\text{}}(X_{i}=x_{i},Y=y)\text{ }\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}$, that is, the set of joint distributions over $(X_{1},X_{2},Y)$ that match the true two-way data distributions. Note that the objective function requires minimization with respect to the three-way mutual information. Importantly,  Bertschinger et al. (2014) demonstrates that the solution to any one objective specifies an optimum solution for all other quantities.
Prior work that relies on this Partial Information Decomposition (PID) to attribute information-theoretic value implicitly assumes that $\Omega_{\text{obs}}^{\text{source}}=\Omega^{\text{source}}=\Omega^{\text{target}}=\Omega$. Instead, we evidence the limitations of these strategies performed on $\Omega_{\text{obs}}^{\text{source}}\neq\Omega$ when the target decomposition is $\Omega^{\text{target}}=\Omega$, i.e., the true data-generating mechanism.
@@
As a motivating example, we analyze the impact of missingness on estimating PID for unidimensional modalities with a bitwise-logic outcome (AND, OR, and XOR). Table 1 (left) presents the discriminative performance associated with neural networks trained on each individual modality and their combination under three scenarios: (i) access to all data (Oracle), (ii) focusing only on datapoints with all covariates observed (Observed), and (iii) adequately accounting for missingness (ICYM2I  using IPW to adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$, by modeling the missingness mechanism), as proposed in Section 4.2. Table 1 (right) presents PID, discussed in Section 4.3 under the same scenarios, demonstrating how information decomposition is also biased due to missingness.
Specifically, relying on $\Omega_{\text{obs}}^{\text{}}$ overestimates the performance of $X_{1}$ for OR but underestimates it for AND. Similarly, biased decomposition overestimates the informativeness of $X_{1}$ (“Unique 1” compared to “Unique 2”) for OR. As $X_{1}$ informs the missingness process, it indirectly informs the outcome of interest, despite the true underlying generative process being dependent on both. The use of IPW can correct for such bias under positivity, provided the IPW propensities can be estimated (i.e., the MAR assumption). We study sensitivity to this assumption in Appendix D, where we further evaluate the robustness of our method under MCAR and MNAR, demonstrating robustness under MCAR.